Grade 12Magical Mathematics[Interesting Approach]

If the binary operation* is defined as a*b=ab+5, then find the value of (a*b)*c, for a=2, b=5 abd c=-2.

shaleen upadhyay
14 Years agoGrade 12

1 Answer

Ashwin Muralidharan IIT Madras
ApprovedApproved Tutor Answer14 Years ago

Hi Shaleen,

 

(2*5) = 10+5 = 15

Now 15*-2 = -30+5 = -25.

-25 is the answer of this binary operation.

 

Hope that helps.

 

Best Regards,

Ashwin (IIT MadraS).